OpenStudy (anonymous):

Solve: 5x − 2(x − 3) = 18 + x A. 6 B. -6 C. 12 D. -12

OpenStudy (ajprincess):

5x-2x+6=18+x 3x+6=18+x 2x=12 x=6

OpenStudy (anonymous):

5x − 2(x − 3) = 18 + x 5x − 2x + 6 = 18 + x 3x − x = 18 + -6 2x=12 x=12/2 x=6 Hope this helps! :D